A brace for a shelf has the shape of a right triangle. Its hypotenuse is 20 inches long and the two legs are equal in length. How long are the legs of the​ triangle?
Answer by Cromlix(4381)

Hi there,
Make the two legs = x
By Pythagoras Theorem
x^2 + x^2 = 20^2
2x^2 = 400
x^2 = 400/2
x^2 = 200
x = √200
x = 14.14 inches (2 decimal places)
Hope this helps :-)